A survey of middle-aged men reveals that 28% of them are balding at the crown of their

heads. Moreover it is known that such men have an 18% probability of suffering a heart

attack in the next ten years. Men who are not balding in this way have an 11% probability of

a heart attack. Find the probability that a middle-aged man will suller a heart attack sometime in the next 10 years


1
Expert's answer
2022-03-17T06:42:58-0400

Let A - "will suller a heart attack", H - "balding". Then, according to the total probability formula

"P(A)=P(H)*P(A|H)+P(\neg H)*P(A|\neg H)=0.28*0.18+0.72*0.11=0.1296"